# Energy, enstrophy and helicity transfers in polymeric turbulence

### v1

@@ 5.2 Local strain
In HIT, $04\left\langle{\alpha\beta\gamma}\right\rangle=-\left\langle{\omega_{i}\omega_{j}S_{ij}}\right\rangle<$ 0 implying $\beta>0$. That is, on an average, we have one large compressive strain is accompanied by two weaker extensional ones. It is known that in HIT, on an average $(\left\langle{\alpha}\right\rangle,\left\langle{\beta}\right\rangle,\left\langle{\gamma}\right\rangle)\approx(3,1,-4)|\beta|$ (Davidson, 2004; Meneveau, 2011). Note that, although this seems to be consistent with the generation of sheet-like structures, it is also consistent with the stretching of vortex tubes, once the self-induced strain is considered (Davidson, 2004).
We now look at the influence of the polymers on the distribution of $\alpha$, $\beta$ and $\gamma$ to relate the $V_{st}(r)$ with the different extension/compression of the fluid elements.
@@
Figure 11: Eigenvalues of $S_{ij}$. (top) Distribution of $s^{*}=-3\sqrt{6}\alpha\beta\gamma/(\alpha^{2}+\beta^{2}+\gamma^{2})^{3/2}$ for different $De$. (bottom) Distribution of (left) $\alpha$, (centre) $\beta$, and (right) $\gamma$ for different $De$.
To this end, we consider the ratio
@@
that was introduced by Lund & Rogers (1994). For a random velocity gradient field with no preferred structure, the distribution of $s^{*}$ is uniform. Note that $s^{*}=1$ is equivalent to $\alpha=\beta=-\gamma/2>0$, and corresponds to a state of axisymmetric extension such that a small spherical fluid element moving in the flow extends (symmetrically) in two directions and contracts in the third. An instance of $s^{*}=0$ corresponds to $\beta=0$, meaning that the state of the straining is two-dimensional and the fluid element is equally stretched/compressed in two directions. In this case, $\alpha\beta\gamma=0$ and the production of enstrophy is null $\omega_{i}\omega_{j}S_{ij}=0$.
The top panel of figure 11 plots the distribution of $s^{*}$ for $0\leq De\leq 9$. As shown by Lund & Rogers (1994), for HIT the distribution peaks at $s^{*}=1$, confirming that the most likely state of straining is an axisymmetric extension. In PHIT, instead, the distribution of $s^{*}$ progressively flattens as $De$ increases, and the most probable value moves towards $s^{*}=0$, i.e. towards a two-dimensional state. The larger preference for $s^{*}=0$ (quasi-two-dimensionality) and not $s^{*}=-1$ is confirmed by the separate distributions of $\alpha$, $\beta$ and $\gamma$ in the bottom panels of figure 11. For the Newtonian case, the results confirm what is found with $s^{*}$, with $\beta$ being mostly positive and $(\left\langle{\alpha}\right\rangle,\left\langle{\beta}\right\rangle,\left\langle{\gamma}\right\rangle)=(4.3,1,-5.3)|\left\langle{\beta}\right\rangle|$. For PHIT, the distribution of the three eigenvalues shrinks and their most likely value monotonically decreases with $De$. To be quantitative, we quote the average values in table 2. This means that in HIT axisymmetric extensions are favoured, while in PHIT shear and planar extensional regions are more frequent. This agrees with the weakening of $V_{st}(r)$ and is consistent with the results by Warwaruk & Ghaemi (2024) that investigated the local flow topology of a polymer-laden turbulent boundary layer. Moreover, these results also validate the suggested mechanism of polymer drag reduction in polymer-laden wall-bounded flows provided for example by Lumley (1973) and Roy et al. (2006). They observed that the extensional viscosity of the polymers strongly inhibits uniaxial and biaxial flow regions, thus favouring two-dimensional states of strain and mitigating the formation of quasi-streamwise vortices.
$De$ $\left\langle{\alpha}\right\rangle/\left\langle{\beta}\right\rangle$ $\left\langle{\gamma}\right\rangle/\left\langle{\beta}\right\rangle$ $\left\langle{\beta}\right\rangle$
$1/9$ $4.3$ $-5.3$ $11$
$1/3$ $4.5$ $-5.5$ $12$
$1$ $5.9$ $-6.9$ $5$
$3$ $7.5$ $-8.5$ $4$
$9$ $9.5$ $-10.5$ $3.5$
@@
Figure 12: (left) $QR$ and (right) $Q_{S}Q_{W}$ maps for (blue) $De=0$ and (red) $De=9$.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 13: Distribution of (left) $Q$ and (right) $R$ for different $De$.
To determine the topology of the flow motion, we focus on two of the invariants of $A_{ij}$, i.e. $Q$ and $R$ (Cantwell, 1993), and plot their distributions in figures 12 and 13. A second-order tensor in three-dimensions possesses three invariants, which are directly related to its eigenvalues $\lambda_{i}$ by means of the characteristic equation
@@
$R=\det(A)=-(\lambda_{1}\lambda_{2}\lambda_{3})=-\alpha\beta\gamma-\frac{\omega_{i}\omega_{j}S_{ij}}{4},$ (35)
@@
where repeated indices are summed over and $tr(A)$ denotes the trace of the generic tensor $A$. The second invariant $Q$ measures the relative strength of strain and vorticity, with $Q\ll 0$ indicating regions of strong strain, and $Q\gg 0$ indicating regions of intense vorticity. The third invariant is a measure of the relative intensity of the production of vorticity ($R<0$) and strain ($R>0$). In figure 12 we also plot the discriminant $\Delta=27R^{2}/4+Q^{3}=0$ ($P=0$) curve for equation 32 with a black line. Below this curve, $\lambda_{i}$s are all real and the flow is dominated by strain. Above this curve, however, $A_{ij}$ has one real and two complex eigenvalues, and enstrophy dominates the flow. When $Q$ is large and positive, strain is locally weak and $R\sim-\omega_{i}\omega_{j}S_{ij}$: in this case $R<0$ implies vortex stretching, while $R>0$ implies vortex compression. When instead $Q$ is large and negative then $R\sim-\alpha\beta\gamma$: a negative $R$ implies a region of axial strain ($\alpha>0$; $\beta,\gamma<0$), while a positive $R$ implies a region of biaxial strain ($\alpha,\beta>0$; $\gamma<0$).
@@
In HIT, the $Q-R$ joint distribution has a tear-drop shape with a clear point at the right-Vieillefosse tail with $\Delta=0$, $R>0$ and $Q<0$ (Ooi et al., 1999; Elsinga & Marusic, 2010). The largest probability is of events lying in the two quadrants where $QR<0$, meaning that there is strong negative correlation between $Q$ and $R$. In other words, the two most common states are vortex stretching $\omega_{i}\omega_{j}S_{ij}>0$ and biaxial strain $\alpha\beta\gamma<0$ (Betchov, 1956; Davidson, 2004). The distributions of $Q$ and $R$ become narrower in PHIT, in agreement with similar observations at small $Re$ by Perlekar et al. (2010). This means that the presence of polymers inhibits the occurrence of vortical and dissipative motions as well as of intense fluid extensions and compressions. This is conveniently visualised in figures 12 and 13 by the shrinking distributions of $R$ with increasing $De$. The $Q-R$ joint distribution for $De=9$ shows that there is still a bias for biaxial extensions in PHIT ($\alpha,\beta,-\gamma,-Q,R>0$). However, stretching is largely diminished compared to HIT, in agreement with more frequent two-dimensional states. The left panel of figure 13 shows that an increase in $De$ leads to a more symmetric distribution of $Q$ showing a stronger inhibition of events with positive $Q$ (large enstrophy). Moreover, the shrinking of the $Q>0$ tail indicates weaker vorticity while the shrinking of the tails of the $R$ distribution hint at a weaker vortical stretching. This is consistent with the discussion in preceding sections.
Figure 14: Qualitative influence of the polymers on the local structure of turbulence. Isocontours of $Q=3Q_{\text{rms}}$ for (top left) $De=0$, (top right) $De=1/3$, (bottom left) $De=1$, and (bottom right) $De=9$ from instantaneous snapshots.
We corroborate these observations with a qualitative picture of the influence of the polymers on the local structure of the flow: see figure 14 where the isocontours of $Q$ are plotted in instantaneous snapshots. For the $De=1$ case, where the fluid-polymer coupling is maximum, the flow features thicker and more elongated vortical structures with more vacant $Q\leq 3Q_{\text{rms}}$ regions. This agrees with the narrower $Q$ distribution discussed above.
@@
$Q_{S}$ and $Q_{W}$ respectively capture the local rates of strain and rotation. These invariants are related to fluid dissipation $\varepsilon_{f}^{\delta q^{2}}$ and fluid enstrophy $\omega^{2}$ as
$\varepsilon_{f}^{\delta q^{2}}=-4\nu Q_{S}\ \text{and}\ \omega^{2}=4Q_{W}.$
@@
Therefore, the $Q_{S}-Q_{W}$ joint distribution also indicates whether the flow is dominated by dissipation or enstrophy. We look at $\mathcal{K}=(-Q_{W}/Q_{S})^{1/2}$ (Truesdell, 1954); when $\mathcal{K}=0$ the flow is extension dominated ($Q_{S}\gg Q_{W})$, when $\mathcal{K}=\infty$ the flow undergoes rigid rotation locally, and is vorticity dominated ($Q_{S}\ll Q_{W}$), when $\mathcal{K}=1$ rotation and stretching are equal, as typical for vortex sheets and shear layers. In HIT, events with $Q_{W}>-Q_{S}$ are more frequent, meaning that the flow is mainly dominated by rigid rotations. In PHIT, instead, the shape of the distribution changes and events with $Q_{W}=-Q_{S}$ ($\mathcal{K}=1$) are favoured. This is in line with the above observation that polymers favour two-dimensional strain states. Note that the influence of polymers on the distribution of $Q_{S}$ ($-\varepsilon_{f}^{\delta q^{2}}$) and $Q_{W}$ ($\omega^{2}$) is rather different. This is visualised in the bottom left panel of figure 13, which considers $Q/Q_{rms}$, with $Q_{S}$ and $Q_{W}$ contributing to the negative and positive values of $Q$, respectively. When $De$ increases, indeed, the positive tail of $Q/Q_{rms}$ ($Q_{W}$) narrows, while the negative one ($Q_{S}$) becomes wider: the presence of polymers suppresses extreme rigid rotation events more than extreme extensions.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 15: Distribution of the cosine of the angle between the vorticity and the eigenvectors of the strain-rate tensor, for different $De$.
As said above, the orientation of $\bm{\omega}$ with respect to the principal axes of $S_{ij}$ is needed to completely determine the local flow topology. We therefore look at the influence of the polymers on the local orientation between the vorticity $\omega$ and $S_{ij}$ (Ashurst et al., 1987; Meneveau, 2011). Figure 15 plots the alignment between the vorticity $\bm{\omega}$ and the principal axes of the strain rate tensor $S_{ij}$. Our results show that polymers favour the anti-alignment between $\bm{\omega}$ and $\hat{\bm{e}}_{\alpha}$ (note that the $\bm{\omega}\cdot\hat{\bm{e}}_{\alpha}=0$ peak monotonically increases with $De$), and favour the alignemnt between $\bm{\omega}$ and $\hat{\bm{e}}_{\beta}$ (note that the $\bm{\omega}\cdot\hat{\bm{e}}_{\beta}=\pm 1$ peak monotonically increases with $De$).
6 Scale helicity
6.1 Scale-by-scale budget
Figure 16: Scale-by-scale helicity budget. In order from top-left to bottom-right, the panels are for $De=0$, $De=1/9$, $De=1/3$, $De=1$, $De=3$ and $De=9$.
We now move to the scale-by-scale helicity budget. In the present set up, turbulence is sustained with the ABC forcing that injects both energy and helicity at the largest scales. As such, like energy, helicity is expected to cascade from large to small scales in the inertial range of scales, and to eventually vanish at the small scales where the mirror symmetry $\bm{u}\cdot\bm{\omega}=0$ is restored.
Our results give evidence of a dual direct cascade of energy and helicity for all $De$ (see figure 16), in agreement with the predictions by Kraichnan (1973), Polifke & Shtilman (1989) and Borue & Orszag (1997) for HIT. For the purely Newtonian case, $\text{d}\langle F^{\delta h}\rangle/\text{d}r\sim(4/3)\langle\varepsilon_{f}^{\delta h}\rangle$ at large scales $r>r_{e}$, $\text{d}\langle\Phi_{f}^{\delta h}\rangle/\text{d}r\sim(4/3)\langle\varepsilon_{f}^{\delta h}\rangle$ in the inertial range of scales $r_{d}<r<r_{e}$, and $\text{d}\langle D_{f}^{\delta h}\rangle/\text{d}r\rightarrow(4/3)\langle\varepsilon_{f}^{\delta h}\rangle$ for $r\rightarrow 0$. Similar to what was observed for $\left\langle{\delta q^{2}}\right\rangle$, polymers provide an additional transfer and dissipative mechanism for helicity, which become more and more important as $De$ increases. Analogous to the cascade of $\left\langle{\delta q^{2}}\right\rangle$, at the intermediate scales the helicity cascade is mainly driven by the nonlinear flux $\langle\Phi_{f}^{\delta h}\rangle$ at scales $r_{p}^{*}<r<r_{e}$ and by the polymeric flux $\langle\Phi_{p}^{\delta h}\rangle$ at scales $r_{d}<r<r_{p}^{*}$. The helicity cascade driven by inertia $\langle\Phi_{f}^{\delta h}\rangle$ is indeed progressively depleted as $De$ increases and the polymers properly stretch, similarly to what was observed for $\langle\Phi_{f}^{\delta q^{2}}\rangle$. Note that the existence of two different cascade processes that take over in two distinct range of scales is consistent with the multiscaling behaviour of $\left\langle{\delta h}\right\rangle$ reported in figure 1.
Refer to captionRefer to captionRefer to caption
Figure 17: Dependence of $\langle\Phi_{f,a}^{\delta h}\rangle$ (left) and $\langle\Phi_{f,b}^{\delta h}\rangle$ (right) on $r$ for different $De$.
Figure 17 provides a closer look at the influence of the polymeric additives on $\langle\Phi_{f}^{\delta h}\rangle$. As mentioned in §2, the nonlinear flux associated with the helicity consists of two distinct contributions. Besides the usual flux term $\langle\Phi_{f,a}^{\delta h}\rangle$ that gives a transfer due to velocity fluctuations $\left\langle{\delta h\delta u_{j}}\right\rangle$, $\langle\Phi_{f}^{\delta h}\rangle$ features an additional term $\langle\Phi_{f,b}^{\delta h}\rangle$ that describes a helicity flux due to $\left\langle{\delta\omega_{j}\delta q^{2}}\right\rangle$. Figure 17 shows that $\langle\Phi_{f,a}^{\delta h}\rangle$ dominates in HIT, being almost twice of $\langle\Phi_{f,b}^{\delta h}\rangle$ at all scales, with helicity being mainly transferred among scales by means of the classic convective term. When increasing $De$, instead, both the contributions to the flux progressively weaken in the elastic range of scales, meaning that polymers deplete both processes while redirecting part of the helicity transfer via the new polymer-driven flux. Notably, polymers have a stronger influence on $\langle\Phi_{f,a}^{\delta h}\rangle$, such that for $De=9$ the two processes on an average contribute equally to the transfer ($\langle\Phi_{f,a}^{\delta h}\rangle\approx\langle\Phi_{f,b}^{\delta h}\rangle$).
6.2 Relative helicity
Refer to captionRefer to captionRefer to caption
Figure 18: (left) Dependence of the relative helicity $r\left\langle{\delta h}\right\rangle/(2\left\langle{\delta q^{2}}\right\rangle)$ on $r$ for different $De$. (right) Dependence of $\left\langle{\delta h}\right\rangle/\left\langle{\delta q^{2}}\right\rangle$ on $r$ for different $De$.
In this subsection, we look at the relative helicity defined as $h_{r}\equiv r\left\langle{\delta h}\right\rangle/(2\left\langle{\delta q^{2}}\right\rangle)$ after Borue & Orszag (1997), to assess how the polymers modify the tendency of the flow to restore mirror symmetry at the smallest scales. Figure 18 shows that the flow is maximally helical at the large scales, and that the relative helicity decays as $r$ decreases. Helicity is indeed dynamically unimportant at small scales. For HIT, we find that $h_{r}\sim r^{1}$ in the inertial range, as also shown by Borue & Orszag (1997). This is visualised in the right panel of figure 18, where $\left\langle{\delta h}\right\rangle/\left\langle{\delta q^{2}}\right\rangle\sim 0.75$, which is similar to the corresponding quantity measured for a passive scalar (see Borue & Orszag, 1997). For PHIT, we find that $h_{r}$ decays at a smaller rate at intermediate $r$ compared to HIT, and that the $r^{1}$ scaling is lost in the inertial range as $\left\langle{\delta q^{2}}\right\rangle$ and $\left\langle{\delta h}\right\rangle$ scale there differently for $De>0$ as already shown in figure 1. We find that $h_{r}$ increases with $De$ at all scales. Overall, in agreement with the distribution of $h=\bm{u}\cdot\bm{\omega}$ in figure 2, the results in figure 18 hint that polymers act to break the mirror symmetry, and that this tendency increases with their elasticity.

### v2

@@ 5.1 Local straining state and alignments
In Newtonian HIT this term is positive, i.e. $\langle\omega_{i}\omega_{j}S_{ij}\rangle>0$, which implies that $\langle\alpha\beta\gamma\rangle<0$. This reflects a characteristic flow configuration in which one large compressive strain is accompanied by two weaker extensional ones. On average, the eigenvalues of $S_{ij}$ follow the approximate ratios $(\langle\alpha\rangle,\langle\beta\rangle,\langle\gamma\rangle)\approx(3,1,-4)|\beta|$ (Davidson, 2004; Meneveau, 2011). Note that, although this seems to be consistent with the generation of sheet-like structures, it is also consistent with the stretching of vortex tubes, once the self-induced strain is considered (Davidson, 2004).
We start looking at the influence of the polymers on the local state of strain of the flow, by analysing the distribution of $\alpha$, $\beta$ and $\gamma$.
@@
Figure 15: Eigenvalues of $S_{ij}$. Distribution of $s^{*}=-3\sqrt{6}\alpha\beta\gamma/(\alpha^{2}+\beta^{2}+\gamma^{2})^{3/2}$ for different DeDe.
To quantify the local strain state, we consider the normalised parameter
@@
originally introduced by Lund & Rogers (1994). For a velocity gradient field with no preferred structure, the distribution of $s^{*}$ is uniform. The limiting values of $s^{*}$ correspond to distinct flow topologies. The value $s^{*}=1$ occurs when $\alpha=\beta=-\gamma/2>0$, representing axisymmetric extension. In this state, a small fluid element symmetrically stretches along two principal directions while contracting along the third. Conversely, $s^{*}=0$ corresponds to $\beta=0$, indicating a two-dimensional straining state where the fluid element experiences equal stretching and compression in two directions. Here, the product $\alpha\beta\gamma=0$ vanishes, and consequently the vortex stretching production term is zero, $v_{st}=\omega_{i}\omega_{j}S_{ij}=0$.
Figure 16: Instantaneous $s^{*}$ field on a two-dimensional slice for $De=0$ (left) and $De=1$ (right). Regions with $s^{*}\leq 0$ (highlighted in blue and white) occur more frequently at $De=1$.
Figure 15 depicts the distribution of $s^{*}$ for $0\leq De\leq 9$. Consistent with Lund & Rogers (1994), the Newtonian HIT case shows a pronounced peak at $s^{*}=1$, indicating that axisymmetric extension is the most probable straining state. In contrast, as $De$ increases in PHIT, the distribution progressively flattens and the mode shifts toward $s^{*}=0$, reflecting a growing tendency toward quasi-two-dimensional strain states. This trend is further supported by direct inspection of the instantaneous flow fields illustrated in figure 16.
DeDe $\left\langle{\alpha}\right\rangle/\left\langle{\beta}\right\rangle$ $\left\langle{\gamma}\right\rangle/\left\langle{\beta}\right\rangle$ $\left\langle{\beta}\right\rangle$
$1/9$ $4.3$ 5.3-5.3 $11$
$1/3$ $4.5$ 5.5-5.5 $12$
11 $5.9$ 6.9-6.9 55
33 $7.5$ 8.5-8.5 44
99 $9.5$ 10.5-10.5 $3.5$
@@
This shift toward two-dimensionality is further supported by the separate eigenvalue distributions of $\alpha$, $\beta$, and $\gamma$ (not shown for brevity). For the Newtonian case, $\beta$ is predominantly positive, with mean values $(\langle\alpha\rangle,\langle\beta\rangle,\langle\gamma\rangle)=(4.3,1,-5.3)|\langle\beta\rangle|$, confirming the dominance of axisymmetric extension. For PHIT, the eigenvalue distributions contract and their modal values decrease monotonically with $De$, indicating a greater prevalence of shear and planar extensional regions (we quote in table 2 the average values).
The increased occurrence of two-dimensional straining explains the observed reduction in the average vortex stretching term $v_{st}$ with increasing $De$ (figure 11). This trend aligns with the results of Warwaruk & Ghaemi (2024), who reported similar modifications in the local flow topology within polymer-laden turbulent boundary layers. Furthermore, these findings support the classic mechanism of polymer drag reduction proposed by Lumley (1973) and Roy et al. (2006), whereby polymer extensional viscosity suppresses uniaxial and biaxial extensional flows, promoting two-dimensional strain states and thereby inhibiting the formation of quasi-streamwise vortices.
Refer to captionRefer to caption
Figure 17: Distribution of the cosine of the angle between the vorticity and the eigenvectors of the strain-rate tensor, for different DeDe.
To further characterise the influence of the polymers on the local flow topology, it is essential to examine the orientation of the vorticity vector $\bm{\omega}$ relative to the principal axes of the strain-rate tensor $S_{ij}$$\hat{\bm{e}}_{\alpha}$, $\hat{\bm{e}}_{\beta}$, and $\hat{\bm{e}}_{\gamma}$ — which correspond to the eigenvectors associated with the eigenvalues $\alpha$, $\beta$ and $\gamma$ (Meneveau, 2011). This alignment is closely linked to the process of vortex stretching, as expressed by the identity (Tsinober, 2001):
$v_{st}=\omega_{i}\omega_{j}S_{ij}=\omega^{2}\alpha\cos^{2}(\bm{\omega},\hat{\bm{e}}_{\alpha})+\omega^{2}\beta\cos^{2}(\bm{\omega},\hat{\bm{e}}_{\beta})+\omega^{2}\gamma\cos^{2}(\bm{\omega},\hat{\bm{e}}_{\gamma}),$ (54)
which decomposes the vortex-stretching term into contributions from each principal strain direction. In Newtonian HIT, the dominant contribution typically arises from the extensional eigenvalue $\alpha$ (Ashurst et al., 1987). Figure 17 shows how this alignment distribution changes with increasing $De$. The results indicate that polymer additives progressively alter the orientation statistics of $\bm{\omega}$. Specifically, as $De$ increases, the probability of $\bm{\omega}$ being orthogonal to $\hat{\bm{e}}_{\alpha}$ increases, as seen from the growing peak at $\bm{\omega}\cdot\hat{\bm{e}}_{\alpha}=0$. At the same time, alignment with the intermediate strain direction $\hat{\bm{e}}_{\beta}$ becomes more pronounced, with peaks at $\bm{\omega}\cdot\hat{\bm{e}}_{\beta}=1$ strengthening with $De$.
These trends suggest that polymers suppress the classical vortex stretching mechanism via two complementary effects: (i) reducing the alignment between vorticity and the most extensional direction, thereby diminishing the dominant $\alpha$-term in $v_{st}$, and (ii) promoting alignment with the intermediate strain direction $\hat{\bm{e}}_{\beta}$, where stretching is weaker, while also increasing the frequency of configurations where $\beta\approx 0$, in which case the associated contribution to vortex stretching vanishes. Together, these modifications lead to the observed reduction in enstrophy production at high $De$.
Figure 18: Terms contributing to $\pi^{\delta\omega^{2}}$ according to equation 55. Left: dependence of the average values of $T^{\omega}||$, $\bm{\nabla}\bm{\omega}||$ and $\rho_{GT}$ on $De$. Right: distribution of $\rho_{GT}$ for different values of $De$.
We now connect the influence of $De$ on $\pi^{\delta\omega^{2}}$, as shown in figure 11, to changes in the flow alignments. Specifically, we write
[…]
Additional insight is provided in the right panel of figure 18, which shows the probability distribution of $\rho_{GT}$ for different values of $De$. In all cases, the distribution peaks near $\rho_{GT}=0$, indicating generally weak alignment. However, for $De<1$, the mode is slightly positive and positive alignment events ($\rho_{GT}>0$) are more probable. For $De\geq 1$, the distribution shifts toward negative values, with negative alignment events ($\rho_{GT}<0$) becoming dominant.
5.2 Invariants of $A_{ij}$
Refer to captionRefer to caption
Figure 19: Top left: The QRQ-R map for $De=0$ (blue) and $De=9$ (red). Top right: The $Q_{S}-Q_{W}$ map for $De=0$ (blue) and $De=9$ (red). Bottom: distributions of $Q$ (left) and $R$ (right) for different DeDe.
To further determine the influence of the polymers on the topology of the flow motion, we now focus on two of the invariants of $A_{ij}$, i.e. $Q$ and $R$ (Cantwell, 1993), and plot their distributions in figure 19. A second-order tensor in three-dimensions $A$ possesses three invariants, which are directly related to its eigenvalues $\lambda_{i}$ by means of the characteristic equation
@@ 5.2 Invariants of $A_{ij}$
$R=\det(A_{ij})=-(\lambda_{1}\lambda_{2}\lambda_{3})=-\alpha\beta\gamma-\frac{\omega_{i}\omega_{j}S_{ij}}{4},$ (59)
@@
where repeated indices are summed over and $tr(A)=A_{ii}$ denotes the trace of the generic tensor $A_{ij}$. The second invariant $Q$ measures the relative strength of strain and vorticity, with $Q\ll 0$ indicating regions of strong strain, and $Q\gg 0$ indicating regions of intense vorticity. The third invariant is a measure of the relative intensity of the production of vorticity ($R<0$) and strain ($R>0$), see Tsinober (2001). In figure 19 we also plot the discriminant $\Delta=27R^{2}/4+Q^{3}=0$ curve for equation (56) with a black line. Below this curve, $\lambda_{i}$s are all real and the flow is dominated by strain. Above this curve, however, $A_{ij}$ has one real and two complex eigenvalues, and enstrophy dominates the flow. When $Q$ is large and positive, strain is locally weak and $R\sim-\omega_{i}\omega_{j}S_{ij}$: in this case $R<0$ implies vortex stretching, while $R>0$ implies vortex compression. When instead $Q$ is large and negative then $R\sim-\alpha\beta\gamma$: a negative $R$ implies a region of axial strain ($\alpha>0$; $\beta,\gamma<0$), while a positive $R$ implies a region of biaxial strain ($\alpha,\beta>0$; $\gamma<0$).
@@
In Newtonian HIT, the QRQ-R joint distribution has a tear-drop shape with a clear point at the right-Vieillefosse tail with $\Delta=0$, $R>0$ and $Q<0$ (Ooi et al., 1999; Elsinga & Marusic, 2010). The largest probability is of events lying in the two quadrants where $QR<0$, meaning that there is strong negative correlation between $Q$ and $R$. In other words, the two most common states are vortex stretching $\omega_{i}\omega_{j}S_{ij}>0$ and biaxial strain $\alpha\beta\gamma<0$ (Betchov, 1956; Davidson, 2004). The distributions of $Q$ and $R$ become narrower in PHIT, in agreement with similar observations at small $Re$ by Perlekar et al. (2010). This means that the presence of polymers inhibits the occurrence of vortical and dissipative motions as well as of intense fluid extensions and compressions. This is conveniently visualised in figure 19 by the shrinking distributions of $R$ with increasing $De$. The QRQ-R joint distribution for $De=9$ shows that there is still a bias for biaxial extensions in PHIT ($\alpha,\beta,-\gamma,-Q,R>0$). However, stretching is largely diminished compared to Newtonian HIT, in agreement with more frequent two-dimensional states. The bottom left panel of figure 19 shows that an increase in $De$ leads to a more symmetric distribution of $Q$ showing a stronger inhibition of events with positive $Q$ (large enstrophy). Moreover, the shrinking of the $Q>0$ tail indicates weaker vorticity while the shrinking of the tails of the $R$ distribution hint at a weaker vortical stretching/compression. This is consistent with the discussion in preceding sections.
@@
$Q_{S}$ and $Q_{W}$ respectively capture the local rates of strain and rotation. These invariants are related to fluid dissipation $\varepsilon_{f}^{\delta q^{2}}$ and fluid enstrophy $\omega^{2}$ as $\varepsilon_{f}^{\delta q^{2}}=-4\nu Q_{S}$ and $\omega^{2}=4Q_{W}$. Therefore, the $Q_{S}-Q_{W}$ joint distribution also indicates whether the flow is dominated by dissipation or enstrophy. We look at $\mathcal{K}=(-Q_{W}/Q_{S})^{1/2}$ (Truesdell, 1954); when $\mathcal{K}=0$ the flow is extension dominated ($Q_{S}\gg Q_{W})$, when $\mathcal{K}=\infty$ the flow undergoes rigid rotation locally, and is vorticity dominated ($Q_{S}\ll Q_{W}$), when $\mathcal{K}=1$ rotation and stretching are equal, as typical for vortex sheets and shear layers. In Newtonian HIT, events with $Q_{W}>-Q_{S}$ are more frequent, meaning that the flow is mainly dominated by rigid rotations. In PHIT, instead, the shape of the distribution changes and events with $Q_{W}=-Q_{S}$ ($\mathcal{K}=1$) are favoured. This is in line with the above observation that polymers favour two-dimensional strain states. Note that the influence of polymers on the distribution of $Q_{S}$ ($\varepsilon_{f}^{\delta q^{2}}$) and $Q_{W}$ ($\omega^{2}$) is rather different. This is visualised in the bottom left panel of figure 19, which considers $Q$, with $Q_{S}$ and $Q_{W}$ contributing to the negative and positive values of $Q$, respectively. As $De$ increases, the positive tail of the $Q$ distribution (associated with rigid-body rotation, $Q_{W}$) narrows more significantly than the negative tail (associated with strain-dominated regions, $Q_{S}$). This indicates that polymers preferentially suppress intense rotational motions over strong extensional events.
Figure 20: Joint conditional average $\langle\Phi_{f}^{\delta q^{2}}(r)|Q_{r},R_{r}\rangle$, for $De=0$ (left) and $De=1$ (right). The scale is set to $r=0.1$, which is within the elastic range of scales. The conditional flux is normalised with $\langle\varepsilon_{f}^{\delta q^{2}}\rangle r$. Positive values of $\Phi_{f}^{\delta q^{2}}$ denote inverse energy cascade, while negative values denote forward energy cascade.
We conclude this section by briefly relating the polymer-induced modifications of local flow topology to their impact on the energy cascade, as discussed in §4.1. Here, we focus on the behaviour of the nonlinear flux term $\Phi_{f}^{\delta q^{2}}$ within the elastic range of scales. Figure 20 displays the joint conditional average $\langle\Phi_{f}^{\delta q^{2}}(r)|Q_{r},R_{r}\rangle$, where
$Q_{r}=\frac{1}{\mathcal{V}(r)}\int_{\mathcal{V}(r)}Q\text{d}\Omega\qquad\text{and}\qquad R_{r}=\frac{1}{\mathcal{V}(r)}\int_{\mathcal{V}(r)}R\text{d}\Omega.$
The scale $r=0.1$ is chosen to be well within the elastic range for sufficiently large $De$. Results are shown for the Newtonian case ($De=0$, left panel) and the viscoelastic case ($De=1$, right panel).
In both cases, $\Phi_{f}^{\delta q^{2}}(r)$ exhibits the characteristic teardrop-shaped structure in the $Q_{r}$$R_{r}$ plane. A pronounced inverse cascade occurs in the quadrant where $Q_{r}>0$ and $R_{r}>0$, corresponding to vortex compression events (Yao et al., 2024a). Conversely, forward cascade activity dominates the quadrants associated with vortex stretching ($Q_{r}>0$, $R_{r}<0$) and biaxial strain ($Q_{r}<0$, $R_{r}>0$), in line with the known role of strain in promoting forward energy transfer (Carbone & Bragg, 2020; Johnson, 2020). While the overall structure remains similar between Newtonian and polymer-laden turbulence, at $De=1$ the region associated with vortex compression and positive $\Phi_{f}^{\delta q^{2}}(r)$ becomes more sharply localised in the normalised $Q_{r}/Q_{r,\mathrm{rms}}$$R_{r}/R_{r,\mathrm{rms}}$ plane. In contrast, the regions linked to vortex stretching and biaxial strain remain more spatially extended. Additionally, whereas the Newtonian case ($De=0$) exhibits the strongest forward cascade intensity in the quadrant with $Q_{r}<0$ and $R_{r}>0$, this peak is less pronounced in the viscoelastic case. In this case, forward cascade intensity becomes more balanced between vortex stretching and biaxial strain events. These findings suggest that, within the elastic range, polymers attenuate the inverse cascade by suppressing vortex compression and weaken the forward cascade mainly through the attenuation of biaxial strain events.